Hi. How would I go about solving this?:
Michico and Patricia are going to play a spinner game. These are the rules:
a) When it is a player's turn, she spins both spinners.
b) Then she adds the two numbers that the arrows point to.
c) If the sum is odd, Michico wins, even if it was not her turn.
d) If the sum is even, Patricia wins, even if it was not her turn.
Patricia tries a test spin first. She spins:
The sum from the first spin is 3, because 3+0=3. Michico wins.
Michico says, "I like this game. I have a better chance to win it than you do."
Patricia says, "No, I have a better chance to win than you do."
Use mathematics to decide which girl is right.
Thanks for your help!

Answer by checkley71(8403)

IF THE NUMBERS ON THE BOARD HAVE THE SAME NUMBER OF ODD & EVEN NUMBERS THEN THE ODDS ARE THE SAME FOR THE SUM OF THE TWO NUMBERS BEING ODD OR EVEN.
0+0=0 EVEN
0+1=1 ODD
0+2=2 EVEN
0+3=3 ODD
0+4=4 EVEN
-
-
-
-
1+0=1 ODD
1+1=2 EVEN
1+2=3 ODD
1+3=4 EVEN

2+0=2 EVEN
2+1=3 ODD
2=2=4 EVEN
2+3=5 ODD